Daisy Farm, a NSW based dairy company exports fresh milk to China. Chinese Yuan (CNY) has been trading at AUD/CNY5.00. Exports to China are currently 100,000 litres per year at the equivalent price of $0.50 per litre. There is a strong rumour that the CNY will be devalued to AUD/CNY$5.60 in the next month. Should the devaluation actually take place, the CNY is expected to remain unchanged for another 10 years. Accepting the rumour/forecast as given, the company faces a pricing decision which must be made before actual devaluation:

The company may either:

  • maintain the same CNY price and in effect sell for fewer dollars, in which case the export volume will not change OR
  • maintain the same dollar price, raise the CNY price in China to compensate for the devaluation and experience a 20% drop in volume. Direct costs in Australia are 60% of the Australian selling price.
  1. What would be the short-run (one-year) implication of each pricing strategy?
  2. Explain which strategy you would recommend to Daisy Farm.
  3. Discuss what other factors Daisy farm should take in to consideration. (Maximum 350 words)
Answer #1

Daisy firm annual milk export to China = 100,000 litres per year

Price per litre = $0.50

Total value of export in Australian dollars = Annual volume * Price per litre

= 100,000 * 0.5

= $50,000

In this case, Price per litre in CNY = Price per litre in AUD * AUD/CNY exchange rate = 0.50 * 5

Price per litre in CNY =CNY 2.50

If Chine Yuan gets devalued from AUD/CNY 5 to AUD/CNY 5.60

1) Under first strategy of maintaining the same CNY price per litre (CNY 2.50) and maintaining the same volume (100,000).

Here price of milk in AUD will get changed due CNY depreciation.

New price in AUD = same CNY price per litre / New AUD/CNY rate after devaluation

=2.50 /5.60

= $0.446

Export Value in AUD = 0.446 * 100,000

Export Value in AUD =$ 44,642.9

Decline in value of export in AUD under first strategy = 44,642.9 - 50,000

Decline in value of export in AUD under first strategy = -$5,357.1

2. Second strategy - maintain same dollar price, increase price in CNY and 20% drop in volume

New volume = 100,000 - 20,000 = 80,000

Value of export in AUD = New volume * same price in AUD

= 80,000 * 0.50

Value of export in AUD = $40,000

Decline in value of export in AUD = 40,000 - 50,000

Decline in value of export in AUD = -$10,000

Since net loss on first strategy of maintaining same price in CNY (-$5,357.1) is lower than second strategy of maintaining same price in AUD (-$10,000)

I will prefer first strategy of maintaining same price in CNY with maintaining the same volumes.

Other factors Daisy firm can consider:

1) If there is rumour about CNY devaluation in the future, Daisy firm can go for hedging in the currency market. They can buy AUD/CHY future contract at forward rate of AUD/CNY 5 so that if CNY gets depreciated against AUD in the future above rate of AUD/CNY 5, future contract will give the positive payoff which will somewhat offset the decline in value of export in AUD.